OpenStudy (anonymous):

what is the least common denominater of 5/12 and -9/16

OpenStudy (welshfella):

one way to do this is the first multiply the denominators 12 * 16 = 192 this is acommon denominator but it might not be the least common denominator

OpenStudy (welshfella):

try dividing it by 2 and see if 12 and 16 are factors of the result

OpenStudy (welshfella):

192 = 96 now is this a common denominator of 12 and 16?

OpenStudy (welshfella):

96 / 12 = ? 96 / 16 = ?

OpenStudy (welshfella):

are the 2 results whole numbers?

OpenStudy (anonymous):

ya

OpenStudy (welshfella):

right so 192 is not the LCD because 96 is a common denominator Now repeat what we did with 96; 96 / 2 = 48 does 12 and 16 divide exactly into 48?

OpenStudy (anonymous):

ya thanks

OpenStudy (welshfella):

well we are not sure if 48 is the LCD yet because we have to try out 48/2 = 24 now is 24 the LCD? does 12 and 16 divide into 24?

OpenStudy (anonymous):

no only 12

OpenStudy (welshfella):

24 /12 = 2 but 16 does not divide into it so we now know that the LCD of 12 and 16 is 48

OpenStudy (anonymous):

We can use prime factorization 12 = 2 x 2 x 3 16= 2 x 2 x 2 x 2 Since they have two pairs of two that overlap we use them each pair once so we get 2 x 2 x 2 x 2 x 3 = 48

OpenStudy (anonymous):

can you help whith one more

OpenStudy (welshfella):

yes that's another method . This method is best for very large numbers.

OpenStudy (welshfella):

please post it separately

OpenStudy (anonymous):

which pair of fractions is equivalent to this pair 3/5 and 2/7 options a. 21/35 and 10/35 B. 10/35 and 7/35 c. 24/35 and 12/35 d. 7/35 and 5/35

OpenStudy (welshfella):

multiply 3/5 and 2/7 by 7/7. Basically what you are doing is multiplying by 1 so the 2 values are retained but the denominator becomes 35